368070:

X Y 1 4.5 2 6.75 3 10.125 4 15.1875 What is the multiplicative rate of change of the exponential function represented in the table? 1.5 2.25 3.0 4.5 can someone please help me answer this

SmokeyBrown:

As X increases, Y also increases, but what can we say about the rate of this change? Well, from 4.5 to 6.75, we can see that 4.5 * 1.5 = 6.75 Then, from 6.75 to 10.125, 6.75 * 1.5 = 10.125 Finally, 10.125 * 1.5 = 15.1875 as well So, from the data, it seems like the multiplicative factor is 1.5

Fan999:

Can I get a screenshot of the question?

368070:

1 attachment
Fan999:

Thank you. Okay so, do you know how to find the rate of change for a function?

368070:

Thank you SmokeyBrown

368070:

thanks fan999 for helping

SmokeyBrown:

No problem!

368070:

Have a good day guys !!

Fan999:

Sure thing :)